OpenStudy (anonymous):

Find all complex numbers z such that z^2=15−8i, and give your answer in the form a+bi. I'm not sure if I have this right

OpenStudy (anonymous):

\[a^2-b^2+2abi=15 -8i\] \[a^2-b^2=15, 2ab=-8\] \[2ab=-8, b=-4a\] \[a^2-(16a^2)=15\] \[-15a^2=15, a^2=-1\] \[a= \pm i\] \[b=-4(i)=-4i\] \[b=-4(-i)=4i\] roots: z= -i+4i=3i z=i-4i=-3i is this right cuz it does not seem right

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

I got to the third line and seen something just a bit off.

OpenStudy (freckles):

2ab=-8 solving this for b gives: b=-4/a not b=-4a

OpenStudy (anonymous):

ooh your right, no wonder I was getting something way off

OpenStudy (anonymous):

oh ok now it makes sense, now it gives me a quartic function and now I can solve for a to get b

OpenStudy (freckles):

yeah did you get: \[a^4-15a^2-16=0 \\ (a^2-16)(a^2+1)=0 ?\]

OpenStudy (anonymous):

yea

OpenStudy (freckles):

cool you will actually get some real values for a which is what you were actually looking for

OpenStudy (anonymous):

Yes thank you :)